OpenStudy (anonymous):

Please can Any One help! A random number generator draws at random with replacement from the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. In 5000 draws, the chance that the digit 0 appears fewer than 495 times is closest to: 25%, 30%, 35%, 40%, 45% or 50%? My solution P(0 appears fewer than 495)=1/10=0.1 n=5000 k<495; K=494 5000C494*(0.1^494) * (0.9^4506)=?

OpenStudy (campbell_st):

makes sense to be... as a binomial probability solution...

OpenStudy (kropot72):

This can be solved by the Normal approximation to the binomial distribution. \[\mu=np=5000\times 0.1=500\] \[\sigma=\sqrt{np(1-p)}=\sqrt{5000\times 0.1\times 0.9}=21.2132\] The z-score for 494 can be calculated, with the continuity correction, as follows: \[z=\frac{494-500+0.5}{21.2132}=-0.25927\] Reference to a standard Normal distribution table shows that the chance that the digit 0 appears fewer than 495 times is closest to 40%.

OpenStudy (anonymous):

Thanks Kropot72

OpenStudy (kropot72):

You're welcome :)